# Yunjue Agent Tech Report: A Fully Reproducible, Zero-Start In-Situ Self-Evolving Agent System for Open-Ended Tasks

### v1

@@ Abstract
Conventional agent systems often struggle in open-ended environments where task distributions continuously drift and external supervision is scarce. Their reliance on static toolsets or offline training lags behind these dynamics, leaving the system’s capability boundaries rigid and unknown. To address this, we propose the In-Situ Self-Evolving paradigm. This approach treats sequential task interactions as a continuous stream of experience, enabling the system to distill short-term execution feedback into long-term, reusable capabilities without access to ground-truth labels. Within this framework, we identify tool evolution as the critical pathway for capability expansion, which provides verifiable, binary feedback signals. Within this framework, we develop Yunjue Agent, a system that iteratively synthesizes, optimizes, and reuses tools to navigate emerging challenges. To optimize evolutionary efficiency, we further introduce a Parallel Batch Evolution strategy. Empirical evaluations across five diverse benchmarks under a zero-start setting demonstrate significant performance gains over proprietary baselines. Additionally, complementary warm-start evaluations confirm that the accumulated general knowledge can be seamlessly transferred to novel domains. Finally, we propose a novel metric to monitor evolution convergence, serving as a function analogous to training loss in conventional optimization. We open-source our codebase, system traces, and evolved tools to facilitate future research in resilient, self-evolving intelligence.
@@ 1 Introduction
The rapid ascendancy of Large Language Models (LLMs) has catalyzed a paradigm shift towards Artificial General Intelligence (AGI) (Guo et al., 2024). While this wave has birthed versatile agents, a critical dichotomy remains: the most capable systems are increasingly dominated by closed-source paradigms (Phan et al., 2025; Gupta et al., 2025). These proprietary models often rely on opaque tools and black-box APIs, creating a barrier for verifiable scientific inquiry. Conversely, open-source alternatives, while accessible, frequently lag in performance (Manchanda et al., 2024; Choi and Chang, 2025), as their reliance on static, manually crafted heuristics renders them unable to rival the robustness of industrial proprietary systems (Jin et al., 2024). We posit that for open-source intelligence to bridge this gap, it must move beyond static imitation of closed models and leverage a dynamic advantage: In-Situ Self-Evolution, where the system becomes stronger through continuous usage.
We envision that true AGI demands an In-Situ Self-Evolving architecture. In this ideal state, an agent acts as a fully adaptive organism, continuously refining its entire cognitive stack—workflow, context and tool. Here, the workflow adapts planning strategies based on user preferences or task types (Ye et al., 2025); the context dynamically distills episodic experiences into wisdom (Tang et al., 2025); and the tools evolve autonomously to conquer novel execution boundaries (Qin et al., 2024).
Recently, the field has witnessed a surge in self-evolving agents (Tao et al., 2024; Gao et al., 2025) designed to enhance these components autonomously (Xi et al., 2025; Cai et al., 2025; Acikgoz et al., 2025). Progress has been made in workflow adaptation (Zhang et al., 2025a) and context management (Tang et al., 2025). However, a critical limitation persists: existing paradigms predominantly rely on offline training with explicit reward signals or remain confined to narrow, domain-specific boundaries.
In supervision-scarce, open-ended environments—where ground truth is absent and user feedback is sparse—this reliance on external supervision becomes a bottleneck. Evolving a workflow or aligning with user preferences often involves subjective or delayed rewards (e.g., “Did the user like this summary?”). In contrast, tool evolution offers a distinct advantage: the feedback signal is intrinsic and objective (Wang et al., 2024a). Regardless of the final task outcome, the functionality of a tool is verifiable via rigorous execution feedback—code either runs successfully or throws an exception. This robust signal allows for autonomous optimization even in the absence of human intervention, making tools the most logical starting point for in-situ evolution. Beyond the signal quality, the toolset also constitutes the fundamental operational prerequisite for any agent (Zaremba et al., 2021). Building on this insight, we advocate for a staged, foundational approach that prioritizes the evolution of tools.
To this end, we introduce Yunjue Agent, a framework that prioritizes the autonomous synthesis, validation, and convergence of tools under a zero-start paradigm. Mechanically, the system employs a multi-agent architecture (Qian et al., 2024), comprising a Manager, Tool Developer, and Executor to address queries dynamically. Instead of relying on a static library, the agent synthesizes bespoke Python primitives on-the-fly when existing capabilities are insufficient. Crucially, to prevent tool explosion and ensure robustness, we introduce a Parallel Batch Evolution strategy. This mechanism processes queries in batches, clustering and merging functionally similar tools via an “absorbing” process. This allows the system to distill transient, query-specific tool into a compact, convergent repository of generalized knowledge.
Our contributions are summarized as follows:
  1. 1.
    SOTA Performance via zero-start In-Situ Self-Evolving: We propose Yunjue Agent, a system that synthesizes, validates, and refines tools from scratch without prior training, with SOTA results on rigorous benchmarks.
  2. 2.
    Cross-domain transferability: We demonstrate that evolved tools exhibit robust transferability across disjoint domains, proving that the system distills transient interactions into generalized, convergent capabilities.
  3. 3.
    Evolution convergence metric: We introduce a quantitative metric to monitor evolutionary stability during inference, serving as a function analogous to “training loss” in conventional optimization.
  4. 4.
    Reproducibility and transparency: We open-source our codebase, complete interaction traces, and the library of evolved tools to provide a transparent foundation for future research.
@@ 2 In-situ self-evolving agents
To enable agents to adapt to new environments like humans, recent work focuses on achieving this through self-evolving methods (Tao et al., 2024; Gao et al., 2025). However, these methods typically require an iterative training process (He et al., 2025; Chen et al., 2025a) to update certain components or are confined in specific areas (Xia et al., 2025; Jin et al., 2025), such as generating query-specific workflows (Zhang et al., 2025a; Ye et al., 2025), refining context $\mathcal{C}$ (e.g., prompt and memory) (Zhang et al., 2025a; Tang et al., 2025) or adjusting toolset (Wang et al., 2024b, 2023). In contrast, we aim to explore whether an agent system can adapt to the environment by continuously updating its components in-situ, where there is no access to external supervision signals. This motivation leads to a new paradigm of agentic evolution.
In-situ self-evolving of agents. Given a sequence of queries $\{x_{1},x_{2},\dots,x_{T}\}$, the agent evolves dynamically from $\mathcal{M}_{0}$ to $\mathcal{M}_{T}$ as it processes each. Specifically, after completing the $t$-th query $x_{t}$, the system updates its configuration for the next query, which can be formalized as a transition from $\mathcal{M}_{t-1}=\langle\mathcal{W}_{t-1},\mathcal{C}_{t-1},\mathcal{T}_{t-1}\rangle$ to $\mathcal{M}_{t}=\langle\mathcal{W}_{t},\mathcal{C}_{t},\mathcal{T}_{t}\rangle$, where any component of the tuple—the workflow structure, context, or toolset—may be modified based on internal feedback or experience gained from the previous interaction. Unlike self-evolving agents that maximize a specific objective function (e.g., accuracy) through a series of training steps (Gao et al., 2025), in-situ self evolving operates during the inference phase where ground truth information is unavailable.
@@ 3.1 In-situ self-evolving via tool accumulation
Upon receiving a query $x_{t}$, the agent first attempts to retrieve relevant utilities from its existing repository, denoted as $\mathcal{T}_{sub}\subseteq\mathcal{T}_{t-1}$. In the absence of requisite capabilities, the agent synthesizes novel tools $\mathcal{P}_{t}$ tailored to the specific task constraints. Subsequently, empowered by the augmented toolset $\mathcal{T}_{sub}\cup\mathcal{P}_{t}$, the agent executes the task through a sequence of tool invocations. To ensure that generated tools are not merely functional but also robust and reusable, the agent engages in a self-reflection mechanism post-execution, refining the tools based on the error reports and execution traces. The process concludes with the integration of these refined tools into the global repository, updating the state to $\mathcal{T}_{t}=\mathcal{P}_{t}\cup\mathcal{T}_{t-1}$.
@@ 3.2 Parallel batch evolution
While sequential query processing maximizes tool reuse and minimizes redundancy, it is inherently inefficient for large-scale benchmarks. Conversely, fully independent parallel processing fails to leverage shared knowledge accumulation across contexts. To reconcile these objectives, we propose a Parallel Batch Evolution strategy that parallelizes agent dynamics while maintaining a cohesive evolutionary trajectory.
Formally, let $\mathcal{Q}_{t}=\{q_{t,1},q_{t,2},\dots,q_{t,B}\}$ denote a batch of $B$ user queries input to the agent $\mathcal{M}_{t-1}$ at step $t$. For each query $q_{t,i}\in\mathcal{Q}_{t}$, the system synthesizes a set of local tools $\mathcal{P}_{t,i}$ that augment the global toolset $\mathcal{T}_{t-1}$ specifically for that instance. However, independent generation often yields functionally redundant tools across different queries—particularly for general-purpose utilities like web searching. This redundancy expands the tool search space, increasing the cognitive load on the agent. To mitigate this, we introduce a tool absorbing mechanism designed to cluster and consolidate utilities post-generation. Specifically, upon completion of batch $\mathcal{Q}$, the system aggregates all tools $\{\mathcal{T}_{t-1},\mathcal{P}_{t,1},\dots,\mathcal{P}_{t,B}\}$ and clusters them into disjoint groups $\{G_{j}\}$ based on functional semantic similarity. Subsequently, a merging function $\Phi$ consolidates these groups, filtering for quality and redundancy to produce a compact updated pool $\mathcal{T}_{t}=\Phi(\{G_{j}\})$. This process ensures the tool space remains streamlined, preventing retrieval ambiguity while updating the system state to $\mathcal{M}_{t}$.
@@ 3.3 The Yunjue agent system
Upon the completion of the batch, the tool absorbing mechanism is invoked to consolidate the dispersed local pools with the global state. Specifically, the system aggregates the union of all toolsets $\{\mathcal{T}_{t-1},\mathcal{P}_{t,1},\mathcal{P}_{t,2},\dots,\mathcal{P}_{t,B}\}$. The LLM-based Aggregator then groups utilities based on functional semantic similarity. Subsequently, another LLM-based Merger is applied to each cluster to synthesize a unified, canonical tool that encapsulates collective capabilities while eliminating redundancy. The resulting consolidated repository $\mathcal{T}_{t}$ serves as the initialized state for the subsequent queries.
@@ 4 Experiment Setup
Datasets. To demonstrate the generalizability of our approach across diverse domains and task complexities, we conduct comprehensive evaluations on four complementary benchmarks, each targeting distinct professional scenarios: (i) HLE (Humanity’s Last Exam) (Phan et al., 2025), a frontier multi-modal benchmark featuring expert-level questions across mathematics, humanities, and natural sciences, designed to assess advanced reasoning at the boundary of human knowledge; (ii) DeepSearchQA (DSQA) (Gupta et al., 2025), which challenges agents’ ability to synthesize comprehensive answers through deep web search, iterative information gathering, and multi-source evidence aggregation; (iii) xBench (Chen et al., 2025b), a Chinese profession-aligned evaluation suite where we focus on the ScienceQA (xSciQA) subset (spanning natural, social, and language sciences) and DeepSearch (xDS, 2025.10 version) subset to evaluate real-world productivity in scientific research and complex retrieval workflows, allowing us to assess cross-lingual adaptation capabilities; and (iv) FinSearchComp (FSC) (Hu et al., 2025), a bilingual (English and Chinese) benchmark targeting financial analysis through its T2 (Simple Historical Lookup) and T3 (Complex Historical Investigation) tasks, which require precise time-sensitive data retrieval and multi-step quantitative reasoning over financial documents.
Baselines. We benchmark our agent system against a comprehensive suite of proprietary and open-source systems, spanning both static and self-evolving agent paradigms. Unless otherwise noted, reported performance metrics are derived directly from original publications, technical reports, or authoritative disclosures (detailed provenance is provided in the Appendix).
@@ 5.1 Zero-start performance on cross-domain tasks
We evaluate our Yunjue Agent, initialized with an empty toolset, across all five benchmarks against state-of-the-art baselines. Results are presented in Figure 2. For datasets excluding xSciQA, foundation model baselines (e.g., GPT-5, Gemini 3 Pro) are augmented with web search and Python interpreters. Conversely, on xSciQA, where standard leaderboards prohibit external tools, we explizcitly equip our backend model with a Python interpreter to ensure a fair parity comparison.
@@
On the heterogeneous HLE benchmark, our agent yields a significant improvement over the backend (48.0 vs. 45.8), ranking second only to GPT-5.2 Pro. Most notably, Yunjue Agent achieves state-of-the-art performance on DSQA, FSC, and xSciQA. We observe substantial absolute gains over the Gemini 3 Pro baseline, with increases of +17.4 points on DSQA (73.5 vs. 56.6) and +15.1 points on FSC (65.0 vs. 49.9), alongside a record-setting score of 76.5 on xSciQA. Finally, on xDS, our method maintains competitive performance (59.7), surpassed only by GPT-5 Pro while outperforming all other models by a significant margin. These exceptional results demonstrate our method’s superior performance across a wide spectrum of tasks, laying a solid foundation for generalized knowledge accumulation and transfer.
@@
Complementing the performance analysis, we aggregated toolsets across all five benchmarks to examine functional utilization patterns. Figure 3 presents the invocation frequency of the top 50 tools. The distribution reveals the spontaneous emergence of high-utility fundamental functions, most notably web_search, fetch_web_content and evaluate_math_expression. The dominance of these foundational tools confirms that the system has effectively distilled generalized knowledge into versatile primitives, ensuring broad applicability across diverse task semantics.
Figure 3: Frequency distribution of the toolset evolved across five benchmarks. We report the top 50 tools, illustrating the emergence of high-generalizability primitives.
5.2 Warm start evolution in shifting domains
We further validate that the generalized knowledge accumulated by our framework is transferable to novel domains, capable of catalyzing capability evolution upon a pre-existing foundation. Specifically, we conduct sequential experiments on DSQA and xSciQA, initializing the system with the toolset derived from the HLE benchmark. HLE is selected as the foundational domain due to its substantial scale (2,500 queries) and extensive disciplinary coverage.
Table 1 presents a comparison of performance and tool synthesis across initialization strategies. We observe that performance remains consistent across domains, with a notable improvement on xSciQA, where the score increases from 76.5 to 80.2. Crucially, the number of newly synthesized tools decreases effectively by 32% on DSQA and remarkably by 100% on xSciQA. This substantial reduction demonstrates that our system’s accumulated generalized knowledge can be seamlessly transferred to disparate domains while retaining the capacity for continuous learning of novel capabilities.
Figure 5: Venn diagram visualization of tool correspondence between zero-start and warm-start settings on DeepSearchQA. The left set comprises tools unique to the zero-start baseline ($\mathcal{T}_{\text{DSQA}}\setminus\mathcal{T}_{\text{HLE}}$), while the right set consists of incrementally generated tools in the warm-start setting ($\mathcal{T}_{\text{HLE}\rightarrow\text{DSQA}}\setminus\mathcal{T}_{\text{HLE}}$). Central intersection indicates high functional overlap. Distinct points represent individual tools, arranged by semantic similarity. Tools in the intersection share similar functionalities, while those within the central dashed lines are exact matches.
Figure 6: Evolutionary generality loss (EGL) on the HLE benchmark and accuracy on a selected dataset as a function of the number of queries. The orange curve represents the EGL trend (left axis). The blue bars show the agent’s accuracy (right axis) at different stages of the evolution (170, 1000, 1750, 2500 queries), highlighting that performance improves and stabilizes as the toolset converges.

### v2

@@ Abstract
Conventional agent systems often struggle in open-ended environments where task distributions continuously drift and external supervision is scarce. Their reliance on static toolsets or offline training renders capability boundaries rigid, preventing adaptation to evolving dynamics. To address this, we propose the In-situ Self-Evolving paradigm. This approach treats sequential task interactions as a continuous stream of experience, enabling the system to distill short-term execution feedback into long-term, reusable capabilities without ground-truth labels. We identify tool evolution as the critical pathway for this capability expansion. Accordingly, we develop the Yunjue Agent, a system that iteratively synthesizes, optimizes, and reuses tools to navigate emerging challenges. To optimize evolutionary efficiency, we introduce a Parallel Batch Evolution strategy. Furthermore, we propose a novel metric to monitor evolutionary convergence, serving as a function analogous to training loss in conventional optimization. Empirical evaluations across five diverse benchmarks under a zero-start setting demonstrate significant performance gains over proprietary baselines and validate the effectiveness of our proposed convergence metric. Additionally, complementary warm-start evaluations confirm that the accumulated general knowledge can be seamlessly transferred to novel domains. We open-source our codebase, system traces, and evolved tools to facilitate future research in resilient, self-evolving intelligence.
@@ 1 Introduction
The rapid advancement of Large Language Models (LLMs) has catalyzed a paradigm shift in the pursuit of Artificial General Intelligence (AGI) (Guo et al., 2024), fundamentally expanding the AI frontier across diverse daily and scientific domains. While these developments have birthed versatile agents, a critical limitation persists: most contemporary systems remain confined within predefined operational envelopes. This reliance on static functional boundaries renders agents inherently fragile when confronted with the stochasticity of open-world environments. We posit that true AGI demands not merely broad-domain competence, but the capacity for self-adaptive evolution—dynamically reshaping capabilities to surmount novel challenges within open-ended contexts. Consequently, architecting a system capable of both cross-domain execution and recursive self-evolution has emerged as a paramount challenge in the quest for resilient intelligence.
Within the framework of LLM-based agents, such adaptive capability typically rests on three pillars: workflow, context, and tool. Recently, the field has witnessed a surge in self-evolving agents (Tao et al., 2024; Gao et al., 2025), designed to autonomously enhance capabilities without human intervention, achieving progress in workflow adaptation (Zhang et al., 2025b), context management (Tang et al., 2025), and tool utilization (Wang et al., 2023). However, existing paradigms often rely on offline supervision (Acikgoz et al., 2025) or remain confined to narrow domains (Wang et al., 2023; Jin et al., 2025). In supervision-scarce, open-ended environments, a more autonomous paradigm is imperative. We argue that among the three pillars of agent systems, while workflows refine expertise and context aligns preference, the toolset constitutes the fundamental prerequisite for novel environments. Without requisite functional tools, even sophisticated planning or memory becomes computationally paralyzed. Furthermore, tool functionality is verifiable via rigorous execution: a program either succeeds or throws an exception. This robust binary signal enables autonomous self-correction even absent human intervention, positioning tools as the optimal starting point for in-situ evolution.
To address the open-ended challenges, we propose the In-situ Self-Evolving (ISE) framework. This paradigm treats task interactions as a continuous stream for accumulating generalized knowledge, initiating with a zero-start phase. We develop the Yunjue Agent, utilizing a multi-agent architecture (Manager, Tool Developer, Executor, etc.) (Qian et al., 2024) to synthesize bespoke Python primitives on-the-fly when existing capabilities are insufficient. To further optimize efficiency, we introduce a Parallel Batch Evolution strategy, which distills high-quality tools from concurrent executions to accelerate convergence. Furthermore, we propose a novel metric to monitor evolutionary convergence, functioning analogously to training loss. We open-source our codebase, complete interaction traces, and the library of evolved tools to provide a transparent foundation for future research in resilient, self-evolving intelligence.
Comprehensive empirical evaluations across five benchmarks validate the Yunjue Agent’s efficacy. Results demonstrate: (i) State-of-the-art performance: In zero-start settings, autonomous tool evolution outperforms proprietary baselines; (ii) Cross-domain transferability: Additional warm-start evaluations show that accumulated general knowledge adapts effectively to novel domains; (iii) Robust convergence monitoring: Our proposed metric effectively serves as a stability proxy for evolution; and (iv) Cost-effectiveness: The lightweight design ensures economical operation.
@@ 2 In-situ self-evolving agents
To enable agents to adapt to new environments like humans, recent work focuses on self-evolving methods (Tao et al., 2024; Gao et al., 2025). However, these methods typically require an iterative training process to update certain components (He et al., 2025; Chen et al., 2025a) or are confined in specific areas (Xia et al., 2025; Jin et al., 2025), such as generating query-specific workflows (Zhang et al., 2025b; Ye et al., 2025), refining context $\mathcal{C}$ (e.g., prompt and memory) (Zhang et al., 2025b; Tang et al., 2025) or adjusting toolset (Wang et al., 2024, 2023). In contrast, we aim to explore whether an agent system can adapt to the environment by continuously updating its components in-situ, where there is no access to external supervision signals. This motivation leads to a new paradigm of agentic evolution.
In-situ self-evolving of an agent system. Given a sequence of queries $\{x_{1},x_{2},\dots,x_{T}\}$, the agent evolves dynamically from $\mathcal{M}_{0}$ to $\mathcal{M}_{T}$ as it processes each. Specifically, after completing the $t$-th query $x_{t}$, the system updates its configuration for the next query, which can be formalized as a transition from $\mathcal{M}_{t-1}=\langle\mathcal{W}_{t-1},\mathcal{C}_{t-1},\mathcal{T}_{t-1}\rangle$ to $\mathcal{M}_{t}=\langle\mathcal{W}_{t},\mathcal{C}_{t},\mathcal{T}_{t}\rangle$, where any component of the tuple—the workflow structure, context, or toolset—may be modified based on internal feedback or experience gained from the previous interaction. Unlike self-evolving agents that maximize a specific objective function through a series of training steps (Gao et al., 2025), in-situ self evolving operates during the inference phase where ground truth information is unavailable.
@@ 3.1 In-situ self-evolving via tool accumulation
Upon receiving a query $x_{t}$, the agent first attempts to retrieve relevant utilities from its existing repository, denoted as $\mathcal{T}_{sub}\subseteq\mathcal{T}_{t-1}$. In the absence of requisite capabilities, the agent synthesizes novel tools $\mathcal{P}_{t}$ tailored to the specific task constraints. Subsequently, empowered by the augmented toolset $\mathcal{T}_{sub}\cup\mathcal{P}_{t}$, the agent executes the task through a sequence of tool invocations. To ensure that generated tools are not merely functional but also robust and reusable, the agent engages in a self-reflection mechanism post-execution, refining the tools based on the error reports and execution traces. The process concludes with the integration of these newly synthesized tools into the global repository, updating the state to $\mathcal{T}_{t}=\mathcal{P}_{t}\cup\mathcal{T}_{t-1}$.
@@ 3.2 Parallel batch evolution
To balance processing efficiency with shared knowledge accumulation, we propose a Parallel Batch Evolution strategy. While sequential processing maximizes reuse, it lacks scalability; conversely, independent parallelization sacrifices shared learning. Our approach reconciles these by parallelizing execution while maintaining a cohesive evolutionary trajectory.
Formally, let $\mathcal{Q}_{t}=\{q_{t,1},q_{t,2},\dots,q_{t,B}\}$ denote a batch of $B$ user queries input to the agent $\mathcal{M}_{t-1}$ at step $t$. For each query $q_{t,i}\in\mathcal{Q}_{t}$, the system synthesizes a set of local tools $\mathcal{P}_{t,i}$ that augments the global toolset $\mathcal{T}_{t-1}$ specifically for that instance. However, independent generation often yields functionally redundant tools across different queries—particularly for general-purpose utilities like web searching. This redundancy expands the tool search space, increasing the cognitive load on the agent. To mitigate this, we introduce a tool absorbing mechanism designed to cluster and consolidate utilities post-generation. Specifically, upon completion of batch $\mathcal{Q}_{t}$, the system aggregates all tools $\{\mathcal{T}_{t-1},\mathcal{P}_{t,1},\dots,\mathcal{P}_{t,B}\}$ and clusters them into disjoint groups $\{G_{j}\}$ based on functional semantic similarity. Subsequently, a merging function $\Phi$ consolidates these groups, filtering for quality and redundancy to produce a compact updated pool $\mathcal{T}_{t}=\Phi(\{G_{j}\})$. This process ensures the tool space remains streamlined, preventing retrieval ambiguity while updating the system state to $\mathcal{M}_{t}$.
@@ 3.3 The Yunjue Agent system
Upon the completion of the batch, the tool absorbing mechanism is invoked to consolidate the dispersed local pools with the global state. Specifically, the system aggregates the union of all toolsets $\{\mathcal{T}_{t-1},\mathcal{P}_{t,1},\mathcal{P}_{t,2},\dots,\mathcal{P}_{t,B}\}$. The LLM-based Aggregator then groups utilities based on functional semantic similarity. Subsequently, another LLM-based Merger is applied to each cluster to synthesize a unified, canonical tool that encapsulates collective capabilities while eliminating redundancy. The resulting consolidated repository $\mathcal{T}_{t}$ serves as the initialized state for the subsequent queries. The pseudo-code of the overall process is provided in Appendix B.
@@ 4 Experiment setup
Datasets. To demonstrate the generalizability of our approach across diverse domains and task complexities, we conduct comprehensive evaluations on five complementary benchmarks, each targeting distinct professional scenarios: (i) HLE (Humanity’s Last Exam) (Phan et al., 2025), a frontier multi-modal benchmark featuring expert-level questions across mathematics, humanities, and natural sciences, designed to assess advanced reasoning at the boundary of human knowledge; (ii) DeepSearchQA (DSQA) (Gupta et al., 2025), which challenges agents’ ability to synthesize comprehensive answers through deep web search, iterative information gathering, and multi-source evidence aggregation; (iii) xBench (Chen et al., 2025b), a Chinese profession-aligned evaluation suite where we focus on the ScienceQA (xSciQA) subset (spanning natural, social, and language sciences) and DeepSearch (xDS, 2025.10 version) subset to evaluate real-world productivity in scientific research and complex retrieval workflows, allowing us to assess cross-lingual adaptation capabilities; and (iv) FinSearchComp (FSC) (Hu et al., 2025), a bilingual (English and Chinese) benchmark targeting financial analysis through its T2 (Simple Historical Lookup) and T3 (Complex Historical Investigation) tasks, which require precise time-sensitive data retrieval and multi-step quantitative reasoning over financial documents.
Baselines. We benchmark our agent system against a comprehensive suite of proprietary and open-source systems, spanning both static and self-evolving agent paradigms. Unless otherwise noted, reported performance metrics are derived directly from original publications, technical reports, or authoritative disclosures (detailed provenance is provided in the Appendix C).
Figure 3: Invocation frequency distribution of the toolset evolved across five benchmarks. We report the top 50 tools, illustrating the emergence of high-generalizability primitives.
@@ 5.1 Zero-start performance on cross-domain tasks
We evaluate our Yunjue Agent, initialized with an empty toolset, across all five benchmarks against state-of-the-art baselines. Results are presented in Figure 2. For datasets excluding xSciQA, foundation model baselines (e.g., GPT-5, Gemini 3 Pro) are augmented with web search and Python interpreters. Conversely, on xSciQA, where standard leaderboards prohibit external tools, we explicitly equip our backend model with a Python interpreter to ensure a fair parity comparison. We provide more detailed results in Appendix D, including the standard deviation of accuracy, token consumption and cost.
@@
On the heterogeneous HLE benchmark, our agent yields a significant improvement over the backend (48.0 vs. 45.8), ranking second only to GPT-5.2 Pro. Most notably, Yunjue Agent achieves state-of-the-art performance on DSQA, FSC, and xSciQA. We observe substantial absolute gains over the Gemini 3 Pro baseline, with increases of +16.9 points on DSQA (73.5 vs. 56.6) and +15.1 points on FSC (65.0 vs. 49.9), alongside a record-setting score of 76.5 on xSciQA. Finally, on xDS, our method maintains competitive performance (59.7), surpassed only by GPT-5 Pro while outperforming all other models by a significant margin. These exceptional results demonstrate our method’s superior performance across a wide spectrum of tasks, laying a solid foundation for generalized knowledge accumulation and transfer.
@@
Complementing the performance analysis, we aggregated toolsets across all five benchmarks to examine functional utilization patterns. Figure 3 presents the invocation frequency of the top 50 tools. The distribution reveals the spontaneous emergence of high-utility fundamental functions, most notably web_search, fetch_text_url and evaluate_expression_math. The dominance of these foundational tools confirms that the system has effectively distilled generalized knowledge into versatile primitives, ensuring broad applicability across diverse task semantics.
Although the Yunjue Agent integrates modules such as the Manager and Tool Developer around the Executor, the Executor acts as the primary entity for execution and token consumption. In Appendix D, we show that across five datasets, the Executor accounts for approximately 95% of the total token consumption. Additionally, since tools do not need to be repeatedly created, a significant amount of token consumption is also saved. Consequently, our system’s overall overhead is comparable to that of a standard ReAct agent equipped with a Python interpreter, a finding further corroborated in Table 2.
5.2 Warm-start evolution in shifting domains
We validate the transferability of accumulated knowledge to novel domains by initializing the system on DSQA, all the way to FSC, xSciQA and xDS, using the toolset derived from the HLE benchmark. HLE is selected as the foundational domain due to its substantial scale (2,500 queries) and extensive disciplinary coverage. Table 1 shows consistent performance across domains, with xSciQA scores improving from 76.5 to 80.2. Crucially, new tool synthesis drops by 55% on FSC and 100% on xSciQA and xDS, demonstrating effective transfer with minimal marginal cost.